# An iterated learning model of language change that mixes supervised and unsupervised learning

### v1

@@ Abstract
The iterated learning model is an agent-based model of language change in which language is transmitted from a tutor to a pupil which itself becomes a tutor to a new pupil, and so on. Languages that are stable, expressive, and compositional arise spontaneously as a consequence of a language transmission bottleneck. Languages are expressive to the extent that a range of different meanings are conveyed unambiguously, and are compositional to the extent that the component parts of meanings are mapped consistently to component parts of the signals that convey them. In the model, a language-using tutor provides a naïve pupil with a limited training set of examples that each pair a meaning with the signal that conveys it in the tutor’s language. The transmission bottleneck ensures that pupils must generalize from the limited training set of meaning-signal pairs. After training, the pupil tutors a new naïve agent, and the cycle repeats. Under the right conditions, the repeated cycles of learning and generalization result in a language that is stable, expressive and compositional. Previous models have implemented an agent’s mapping from signals to meanings using an artificial neural network decoder, but have relied on an unrealistic and computationally expensive process of obversion to implement the associated encoder, mapping from meanings to signals. Here, a new model is presented in which both decoder and encoder are neural networks, trained separately through supervised learning, and trained together through unsupervised learning in the form of an autoencoder. This avoids the substantial computational burden entailed in obversion and introduces a mixture of supervised and unsupervised learning observed during human development. The model achieves more complex expressive compositional languages than previous studies, and reveals a linear relationship between the dimensionality of a language’s meaning-signal space and the size of the optimal transmission bottleneck.
@@ Introduction
Languages are made easier to learn and easier to interpret by their expressivity, compositionality and stability; that is, by the tendency for a language to be able to express many different meanings using different sentences, to compose meaningful sentences from meaningful sub-parts and to change slowly over generations. There are different words for left and right, for example, but the word ‘left’ in ‘you write with your left hand’ means something very similar to the word ‘left’ in ‘turn left at the junction’. Moreover, the difference in meaning between the expressions ‘turn left, then right’ and ‘turn right, then left’ is represented by the different structural organization of their component words. Of course, languages are neither completely expressive nor completely compositional, ‘they are on the left’ typically refers to position but, in some contexts, could refer to politics, and a sailor in some cases would use the word ‘port’ when left might be expected. Nonetheless, expressivity and compositionality are, invariably, important features of languages and might be considered to mark the boundary between a simple system of signs and utterances and a true language. The third property, stability, is also important. Since parents need to be able to understand what their children say to them, at least some of the time, it is important that language does not change drastically from generation to generation. Obviously these are only three of a long list of properties a ‘good’ language needs in order that it be easy to learn and effective for communication; however, these three are among the most basic and are our focus here.
@@
The iterated learning model (ILM) is an agent-based model of language change that demonstrates how a transmission bottleneck can encourage the spontaneous emergence of a language that is expressive, compositional and stable. Here a language is defined as a map from a set of ‘meanings’ to a set of ‘signals’ and, broadly, the goal for the iterated learning model is to shed light on the conditions under which languages that are expressive, compositional and stable arise. We will define measures, $x$, $c$ and $s$ for expressivity, compositionality and stability; these measures all take values in $[0,1]$ and we define an XCS language as one where all three exceed $\lambda=0.95$. In our limited context this will serve as a definition of a ‘good’ language.

### v3

@@ Abstract
The iterated learning model is an agent model which simulates the transmission of of language from generation to generation. It is used to study how the language adapts to pressures imposed by transmission. In each iteration, a language tutor exposes a naïve pupil to a limited training set of utterances, each pairing a random meaning with the signal that conveys it. Then the pupil becomes a tutor for a new naïve pupil in the next iteration. The transmission bottleneck ensures that tutors must generalize beyond the training set that they experienced. Repeated cycles of learning and generalization can result in a language that is expressive, compositional and stable. Previously, the agents in the iterated learning model mapped signals to meanings using an artificial neural network but relied on an unrealistic and computationally expensive process of obversion to map meanings to signals. Here, both maps are neural networks, trained separately through supervised learning and together through unsupervised learning in the form of an autoencoder. This avoids the computational burden entailed in obversion and introduces a mixture of supervised and unsupervised learning as observed during language learning in children. The new model demonstrates a linear relationship between the dimensionality of meaning-signal space and effective bottleneck size and suggests that internal reflection on potential utterances is important in language learning and evolution.
Author summary
Languages tend to be expressive: distinct sentences communicate distinct meanings, and compositional: sentence parts consistently correspond to meaning parts. In the iterated learning model a tutor teaches a pupil their language, in this case a mapping from binary vector ‘meanings’ to ‘signals’, by exposing them to a set of meaning-signal pairs. Subsequently, the pupil becomes the tutor for a new naïve pupil. Successive cycles of learning and generalization lead to an expressive, compositional and stable language. However, while an agent’s map from signals to meanings is implemented by its neural network, the complementary map from meanings to signals is calculated through ‘obversion’ which becomes impractical for larger languages and is unrealistic, requiring consideration of all possible meaning-signal pairs. Here, we propose a more computationally tractable and more realistic model in which both the encoder and decoder are neural networks, concatenated together as an autoencoder, and pupils are required to learn from a mix of unsupervised and supervised examples, as children do. The model demonstrates the emergence of languages that are expressive, compositional and stable, and reveals a linear relationship between the optimal size of the transmission bottleneck and the dimensionality of the language’s meaning-signal space.
@@ Introduction
In this paper we introduce a new iterated learning model (ILM) for language learning; compared to the existing model, this new model is less computationally demanding and is closer to human language learning.
One of the most important questions in linguistics asks to what degree our linguistic ability prepares us for the specific forms that human languages take [1, 2, 3]. It may, however, be possible to sidestep this question by asking two perhaps more pertinent questions: (1) what specific properties do human languages have that make it possible for us to to learn and interpret them and to use them to communicate about the diverse and often unforeseen topics we need to discuss and (2) how do languages come to have these properties? Languages are made easier to learn and easier to interpret by their expressivity, compositionality and stability; that is, by the tendency for a language to be able to express many different meanings using different sentences, to compose meaningful sentences from meaningful sub-parts and to change slowly over generations. There are different words for left and right, for example, but the word ‘left’ in ‘you write with your left hand’ means something very similar to the word ‘left’ in ‘turn left at the junction’. Moreover, the difference in meaning between the expressions ‘turn left, then right’ and ‘turn right, then left’ is represented by the different structural organization of their component words. Of course, languages are neither completely expressive nor completely compositional, ‘they are on the left’ typically refers to position but, in some contexts, could refer to politics, and a sailor in some cases would use the word ‘port’ when left might be expected. Nonetheless, expressivity and compositionality are, invariably, important features of languages and might be considered to mark the boundary between a simple system of signs and utterances and a true language [4, 5, 6]. The third property, stability, is also important. Since parents need to be able to understand what their children say to them, at least some of the time, it is important that language does not change drastically from generation to generation.
Expressivity, compositionality and stability are only three of a long list of properties a ‘good’ language needs in order that it be easy to learn and effective for communication. However, these three are among the most basic, indeed compositionality is a version of Frege’s principle of compositionality [7, 4] which is central to his very influential theory of language. Here, we focus on expressivity, compositionality and stability because of their importance and because they can be measured in the sort of simple language used in our simulations. In ILM-based investigations of how language can acquire important basic properties through iterated learning, these three are intended to stand in for the multiple properties needed for a language to be useful and learnable.
We look at the ILM, [8, 9, 10, 11], a model of learning in which a language evolves across generations; this model provides a putative description of how a language might become expressive, compositional and stable. However, as we describe below, the model, requires an ‘obversion’ step which makes it computationally expensive and unrealistic as a simulation of human language learning. Here we propose a solution to this and argue that our new model suggests the importance to language learning and to language evolution of an internal process within the language learner, translating from some meaningful state or idea to an utterance that could be used to express it, and back to how that the meaning of that putative utterance might be interpreted.
@@
The ILM models how a language evolves as a tutor teaches their language to a pupil and the pupil, in turn, becomes a tutor and teaches their language to a new pupil. The model demonstrates the gradual emergence of a language that is expressive, compositional and stable, despite the fact that at every iteration the pupil begins in an identical state of ignorance. Here a language is defined as a map from a set of ‘meanings’ to a set of ‘signals’ and, broadly, the goal for the iterated learning model is to shed light on the conditions under which languages that are expressive, compositional and stable arise. We will define measures, $x$, $c$ and $s$ for expressivity, compositionality and stability; these measures all take values in $[0,1]$ and we define an XCS language as one where all three exceed $\lambda=0.95$. In our limited context this will serve as a definition of a ‘good’ language; their precise definitions are giving in S1 Appendix..
The ILM was introduced in [8, 9, 10]. In its original form, it demonstrated that the emergence of language compositionality can be attributed to the influence of a transmission bottleneck on language acquisition taking place over a number of successive generations [9, 12]. This original version of the model was used to explain the mixture of irregularity and regularity found in languages [9]. However, this ILM used a language generated by a simple formal grammar and relied on a set of deduction rules that the pupil used to deduce the formal grammar from exemplars. This made it difficult to disentangle the role of the deduction rules from the iterated dynamics of language learning. This was addressed in [11] where a new version of the ILM is presented which did not require these deduction rules.
This ILM has a decoder based on a neural network and incorporated obversion, an algorithm to derive an encoding map from the decoding map, which had been previously defined in a more general context in [13]. This obversion step is problematic in two ways: it is computationally expensive, limiting the ILM to small example languages and it is unrealistic, it makes the transition from pupil to tutor a discrete step during which the pupil must calculate the probability associated with each signal as an encoding of every possible meaning, something very different to what happens in human language learning. Our purpose here is to propose an ILM which does not require obversion.
This is not to suggest that the Obverter ILM has not been useful. For example, since it was introduced an obverter-based learning models has been applied to pictures [14]. This is exciting since it simulates the communication aspect of language use and links this class of language-learning model to classic Lewis signalling games [15]. It is hoped that an obversion-free version based on our proposal here would make it computationally possible to investigate this further. Population-based versions of the ILM have been considered [16, 17, 18], this is important since language evolution involves space and even geography as well as time, but the Obverter ILM has the disadvantage that an agent cannot learn and teach at the same time, something the new model introduced here will address.